OpenStudy (anonymous):

i nedd a correct answer for Mr. Alvirez has 100 feet of fencing. What are th diemensions of the largest rectangular pen he can build for his dogg

OpenStudy (angela210793):

10x10

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Let \(x\) and \(y\) be the length and the width of the pen, respectively. We have \(2x+2y=100 \implies x+y=50\). We want to maximize the area, which is \(xy=x(50-x)=50x-x^2\). The maximum value occurs at \(50-2x=0 \implies x=25\). Sot he dimensions of the largest pen are \(25\times25\).

OpenStudy (anonymous):

\(10\times10\) would be true if Mr. Alvirez has only 40 feet of fencing.

OpenStudy (anonymous):

anwara@ does yours equal too 100

OpenStudy (anonymous):

Yeah, The correct dimensions are 25x25.

OpenStudy (anonymous):

so its not 10x10

OpenStudy (anonymous):

It's NOT correct!